Staff:

LBTO:

Josh, Juan Carlos, Andrew

Astronomer: remote

ALTA.

alta0_20221108.png alta1_20221108.png

Summary

Details.

Telescope setup for Luci+AO

4:30 PM Power on

Ocam
  • not live.
  • masterdiagnostic:
  • UNEXPECTED BCU FRAME size.
  • bcu2k power cycle.
  • reset the bin 1x1 cropped.
  • OK
ocam_notlive_20221108.png

OSCO

5:42 POINTING PRESET.

ELEVATION 56

seeing guider 1.1, variable +-0.2"

sky looks clear.

guider-20221108_0.png

some images of the guider 0.9"

00:55 starting pointing.
  • flux mag: 8.5
  • 0ptimize: 20221109_005954:
  • seeing jumps to 1.2"
  • first pass of the TT ->adsec failure. (??)
  • idlctrl.L.202211090000.log:idl_ctrl_19010 |INF| 153776|2022-11-09 01:00:23.640789| MAIN > Executing command: print, fsm_set_gain('/local/aomeas/adsec_calib/M2C/KL_v20/gain/g_tt_22.38_ho1_0.10_ho2_0.10.fits')

    idlctrl.L.202211090000.log:idl_ctrl_19010 |INF| 153781|2022-11-09 01:00:23.722000| MAIN > Executing command: print, fsm_set_gain('/local/aomeas/adsec_calib/M2C/KL_v20/gain/g_tt_22.88_ho1_0.10_ho2_0.10.fits')

    idlctrl.L.202211090000.log:idl_ctrl_19010 |INF| 153786|2022-11-09 01:00:23.800155| MAIN > Executing command: print, fsm_set_gain('/local/aomeas/adsec_calib/M2C/KL_v20/gain/g_tt_23.38_ho1_0.10_ho2_0.10.fits')

    idlctrl.L.202211090000.log:idl_ctrl_19010 |INF| 153791|2022-11-09 01:00:23.889123| MAIN > Executing command: print, fsm_set_gain('/local/aomeas/adsec_calib/M2C/KL_v20/gain/g_tt_23.88_ho1_0.10_ho2_0.10.fits')

  • The TT gain too high..
6:03 PM adsc_stop/start
  • elevation 63
  • flat 641 modes.
  • repeat preset.
6:06PM
  • center pupil failed.
  • the camera lens was in positions, but the process seems not moving the camera lens and keeping the same error in the pupils firts was x=-2 y=-1
  • moving the cl by hand the error reduced to 0.5 and the pupil centering didn;t reduce the error.
  • restarted the process pupil check ocam
  • no luck.
restart WFS soft
  • bcu47 not reset, tried two times and it didn; excuted the reset, lost comm with the bcu47.
restart w soft.
  • ocam
  • ok
  • Issue with the DIMM connectivity
  • contacted the soft support.
  • Science preset.
  • 6:41PM
  • check camera lens move manually.
  • science target.
  • The Preset AO resports the pupil rerotator out of range.
  • pupil rerotator pos:349.588

pupil rerotator position should be at the time to prepared the OT. How to pass this info.. (??)

6:56PM
  • wind speed peak to 4m/s
  • Preset AO
  • AO preset@01:58:55.885 with a modified rot angle by 90 degrees
  • it seems the LFBG rotator didn;t move.
  • rotator moved after a few seconds idle, the WFS rerotarot moved and it shows tracking positions enabled. The positions 335.
  • IIF reported back an error that the pupil rerotator was out of range (??).
  • 01:59:21.417 WFS rerotator out of range.
  • It seems a timing of updating variables and status of the RTDB
  • IT open.IT8762
  • Target too low ot be observed and wind ramping up, gust insdie dome 5m/s.
  • No other targets for AO
Change to seeing limited.

WFS power off

10:01PM

-- JuanCarlosGuerra - 08 Nov 2022
Topic revision: r2 - 09 Nov 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback